OpenStudy (alexis_t.ortiz.):

Write a recrusive formula for the sequence 5, 18, 31, 44, 57 Then find the next term A) a1=-8;an=an-1+13 B)a1=5;an=an-1+13 C)a1=5;an=an-1-13 D)a1=-8;an=an-1-13

OpenStudy (jango_in_dtown):

18-5=31-18=44-31=..=13

OpenStudy (jango_in_dtown):

also first term is 5

OpenStudy (jango_in_dtown):

its an ap series

OpenStudy (jango_in_dtown):

with common difference 13

OpenStudy (welshfella):

basically each term after 5 is obtained by adding 13 to the previous term.

OpenStudy (jango_in_dtown):

enough hints for the solution I guess

OpenStudy (alexis_t.ortiz.):

Thank you!